OpenStudy (anonymous):

What are the roots of 2x^2-3x+1

OpenStudy (anonymous):

luckily, u can factor this one. so: (2x-1)(x-1)

OpenStudy (anonymous):

Some1 is correct. The roots can be found by setting 2x-1=0 So x=1/2 and by setting x-1=0 so x=1